### v4

# Unveiling Molecular Moieties through Hierarchical Grad-CAM Graph Explainability

Abstract
Background: Virtual Screening (VS) has become an essential tool in drug discovery, enabling the rapid and cost-effective identification of potential bioactive molecules. Among recent advancements, Graph Neural Networks (GNNs) have gained prominence for their ability to model complex molecular structures using graph-based representations. However, the integration of explainable methods to elucidate the specific contributions of molecular substructures to biological activity remains a significant challenge. This limitation hampers both the interpretability of predictive models and the rational design of novel therapeutics.
Results: We trained 20 GNN models on a dataset of small molecules with the goal of predicting their activity on 20 distinct protein targets from the Kinase family. These classifiers achieved state-of-the-art performance in virtual screening tasks, demonstrating high accuracy and robustness on different targets. Building upon these models, we implemented the Hierarchical Grad-CAM graph Explainer (HGE) framework, enabling an in-depth analysis of the molecular moieties driving protein-ligand binding stabilization. HGE exploits Grad-CAM explanations at the atom, ring, and whole-molecule levels, leveraging the message-passing mechanism to highlight the most relevant chemical moieties. Validation against experimental data from the literature confirmed the ability of the explainer to recognize a molecular pattern of drugs and correctly annotate them to the known target. Conclusion: Our approach may represent a valid support to shorten both the screening and the hit discovery process. Detailed knowledge of the molecular substructures that play a role in the binding process can help the computational chemist to gain insights into the structure optimization, as well as in drug repurposing tasks.
footnotetext: corrisponding author salvatore.contino01@unipa.it (Salvatore Contino)footnotetext: 1Department of Engineering, University of Palermo, 90128 Palermo, Italyfootnotetext: 2Molecular Informatics Group,Fondazione Ri.MED, 90128 Palermo, Italy
1 Background
The development of a new drug is a long and time-consuming process that, despite advances in technology and computer applications, is fragmented into several steps toward the identification of a lead compound from in vivo models. During the last decade, the use of deep learning has provided support to Virtual Screening (VS) campaigns as an effective predictive means to select hit compounds through the use of massive data in the training model phase [16, 5, 31, 3, 40]. In fact, being able to correctly prioritize active molecules during the screening steps speeds up the discovery process. The classical approaches used for virtual screening are not always efficient, being affected, for example, by the use of specific force fields [30, 20] or the quality of structural data available, thus affecting the whole discovery process from hit identification to hit optimization steps towards the lead compound. Hit discovery is the very first step of a drug discovery workflow, and chemists use it as the starting point to rationally optimize molecules up to the preclinical candidate. The molecular optimization steps involve different techniques such as hit expansion [2], (bio)isosteric replacement [9] or a combination of them [15]. Thus, optimization requires an understanding of the relevant chemical features within the small molecule. This step is usually handled by computational chemists who recognize the molecule’s activity and consciously act to improve it in accordance with their knowledge and expertise in the field.
In this context, Graph Neural Networks (GNNs) [28, 17] have emerged as a powerful tool in computational drug discovery, offering unique capabilities to model molecular structures [4, 34, 33, 14]. Unlike traditional machine learning models that rely on precomputed molecular descriptors, GNNs operate directly on graph representations of molecules, where atoms are nodes and bonds are edges [13]. This allows GNNs to effectively capture the intrinsic connectivity and topology of molecular structures.
Through iterative message passing mechanisms [10], GNNs propagate information across the graph, thus enabling the model to learn complex relationships between the atom neighborhoods and the molecular properties. This data-driven approach reduces the dependence on hand-crafted features, instead leveraging the model’s ability to extract relevant patterns directly from the molecular graph. In addition, GNN architectures can be tailored to specific tasks, such as predicting molecular properties, bioactivity, or drug-target interactions by integrating domain-specific information through customized layers and loss functions [39, 35, 6, 32].
Even if the graph representation of a molecule is self-explainable, the precise relation between the relevant molecular properties and the network prediction is not, and it suffers from the same issues observed in classical Convolutional Neural Networks (CNN) devoted to images. As a consequence, graph-oriented eXplainable Artificial Intelligence (XAI) techniques have emerged in the Drug Discovery domain. A very recent work by Proietti et al. [24] addresses the Concept Withening (CW) technique to investigate the decision process of a Graph Convolutional Neural Network (GCNN) aimed at discovering relevant molecular properties on a plethora of benchmark data sets. CW has been presented in [7]: it is a representation learning technique aimed at aligning the axes in the latent space close to some known “concepts” (i.e. group of features) in the data space. Proietti and his colleagues used GNNExplainer [36] to learn the contributions of concepts in the predictions of a QSAR model. GNNExplainer is an explainer capable of identifying compact subgraphs and small subsets of features through the use of Mutual Information ($MI$) allowing the GNNExplainer to be formulated as:
(1) $\max_{G_{S}}MI\left(Y,\left(G_{S},X_{S}\right)\right)=H(Y)-H\left(Y\mid G=G_{S},X=X_{S}\right)$
The $MI$ term measures how much information about the predicted label distribution $Y$ is gained by knowing both a suitable subgraph $G_{S}$ and a subset of the features $X_{S}$. By maximizing $MI$, the model seeks the subgraph $G_{S}$ and the features subset $X_{S}$ that provide the most information about $Y$. This is achieved by minimizing the conditional entropy $H\left(Y\mid G=G_{S},X=X_{S}\right)$, which indicates that knowing the chosen subgraph and features reduces the uncertainty in predicting $Y$ as much as possible. In addition to the GNNExplainer, other XAI frameworks have been adapted to explain the information obtained from GNNs. Specifically, the two most widely used methods are SHapley Additive exPlanation (SHAP) and Gradient-weighted Class Activation Mapping (Grad-CAM). SHAP aims to explain the contribution of individual nodes and edges (features) in a GNN to the final prediction for a molecule. To assess the importance of either node or an edge, SHAP creates a “coalition” centered around that specific feature. This coalition includes the node/edge itself and all its descendant nodes and edges in the graph. SHAP then compares the prediction of the model for the molecule with and without the entire coalition. This difference reflects the impact of the target node/edge and its connected features on the prediction. By analyzing SHAP values for different coalitions centered on individual nodes and edges, researchers gain insight into the relative importance of each feature for the GNN’s prediction. In the equation 2, $g$ is the explanation model, $z^{\prime}\in{0,1}^{M}$ is the coalition vector, $M$ is the maximum coalition size and $\phi_{j}\in\mathbb{R}$ is the $j$-th feature attribute.
(2) $g\left(z^{\prime}\right)=\phi_{0}+\sum_{j=1}^{M}\phi_{j}z_{j}^{\prime}$
The equation shows how the SHAP framework breaks down model predictions into contributions from single features or feature combinations, giving each a weight determined by its relative importance. In [26, 25], SHAP was used to determine the significance of individual features; however, it did not provide insight into their pharmacophoric characteristics.
On the other hand, Grad-CAM [29] is an explainability technique widely used in image classification through the use of CNNs. It generates a heat map for each of the classes to highlight the parts of an image that contribute to the recognition of that class. At each layer where Grad-CAM is applied, the role of each feature map in explaining the class is weighted using the gradients backpropagated to the layer itself. In [23], a graph-based adaptation of Grad-CAM is introduced, which is called GCNN-Explainability. This method incorporates a Grad-CAM layer in the final stage of the GCNN architecture to identify substructures, that play a significant role in individual classification, thus highlighting features that are representative of a specific class. The application of the Grad-CAM technique to graph-based models, particularly in the context of molecular graphs, offers significant advantages. By visualizing the most relevant graph components for a classification task, these methods can shed light on the key pharmacophoric elements within a molecule. This not only aids in understanding the decision-making process of the model, but also provides valuable insights for drug discovery and molecular design, by focusing attention on the structural components that drive biological activity.
In this paper, we present the Hierarchical Grad-CAM graph Explainer (HGE) framework, which is designed to identify the most relevant molecular substructures, that influence activity prediction in two data sets of small molecules on twenty different Kinase protein targets. HGE leverages Grad-CAM to obtain local explanations at atom, ring, and the entire molecule level, and then it combines local explainers to provide information on the most relevant molecular moieties. To prove the effectiveness of HGE we built twenty GCNN classifiers, each dedicated to a specific protein target. These models were trained to optimize classification performance on a single target. In each classification task, the ground-truth activity labels were sourced from ChEMBL. The proposed HGE framework extends Grad-CAM by providing a deeper analysis of the relevant sub-graphs directly linked to the chemical functionality of the identified moieties, focusing on the substructures directly involved in predicting an active molecule, rather than providing a global explanation of the classifier behavior. These substructures or portions of them are expected to play a crucial role in binding to the target. As a consequence, we used a hierarchy of explainers placed at different layers to catch information directly at the atom, ring, and whole molecule level by leveraging the message passing mechanism that is used in GNNs.
The analysis carried out on a set of already approved drugs demonstrates that trained neural networks can successfully identify inhibitors for each of the 20 targets with a high success rate. HGE proved to be very effective in detecting the same common substructures in different molecules that are active on the same target. Conversely, HGE selected the proper diverse substructures for the same molecule when its activity is investigated versus different targets.
Next sections of the paper are arranged as follows: Section 2 describes the creation of the two data sets used for training the networks, their structure, and the development of the explainability algorithm. Section LABEL:sec3 reports the classification performance analysis of the implemented GNNs with respect to different metrics, and the results obtained from HGE, along with a discussion of the interpretability of these results. Finally, Section LABEL:sec5 sets the groundwork for future studies.
2 Methods
All the experiments were carried out using two data sets: EMBER and ChEMBL_over. EMBER was curated by some of the authors [21], and it was already used in previous research. Data were appropriately encoded in the graph format for the purposes of this work. ChEMBL_over has been designed purposely for this study with the aim of stressing the classification performance in case of extreme unbalanced data. The data set was derived from the ChEMBL database [38] and enriched by oversampling inactive compounds for each target. The following sections will provide a detailed explanation of the creation processes for both data sets.
2.1 Data preparation
2.1.1 Data set 1: EMBER
The first data set used in this paper has been proposed by some of the authors in the work by Mendolia et al. [21]. The input molecular graphs used by our GCNNs were generated from the canonical SMILES representation. Unlike the other molecular descriptors used in classical approaches for Virtual Screening, molecular graphs account for chemical information both in each atom and its chemical neighborhood. Canonical SMILES were recovered from the CHEMBL24.1 database [38] looking for all small molecules whose activity on the target protein is referenced by the values of IC50, Ki, and Kd. We used a fixed threshold activity value on each target in order to generate the molecular graphs through the use of the RDKit library.
Each node represents an atom as a feature vector and nine features were selected in the proposed representation:
  • nosep
    Atomic number
  • nosep
    DegreeAn atom’s degree is defined as the number of directly bound neighbors. The degree is independent of bond orders, however it depends [18]
  • nosep
    Formal Charge
  • nosep
    Hybridization
  • nosep
    Aromaticity
  • nosep
    Total number of hydrogen atoms
  • nosep
    Number of radical electrons
  • nosep
    Information on aromaticity
  • nosep
    Chirality
The set of atom features was later augmented with the 3D coordinates of each atom in active configuration, for the sake of training the single-class GCNN architecture used for the hierarchical explainability analysis with respect to each Kinase protein target. These coordinates were calculated by minimizing the conformation of the molecule and creating the correct protonation state using Ligprep [1] with $pH=7.4$. In each molecular graph, the edge information encodes the type of bond (double or triple bond) and whether the bond is within a cycle. The size of each molecular graph had several nodes and arcs which vary according to the size of the molecule.
As reported in the paper by Mendolia et. al, inactive compounds in the EMBER data set were devised using the opposite of the concept of similarity, that is, dissimilarity. A compound is labeled as inactive when it is “dissimilar” to the target: its similarity coefficient to the active compounds is less than $0.1$. At first, several inactive compounds were discovered with respect to all the targets on ChEMBL. Then each class was further enriched with compounds that are inactive specifically on the target even if they are active on other targets. The final data set consists of 89373 small molecules and has a 1:100 active/inactive ratio for the smallest class, which refers to the CLK2 target protein. This ratio allowed us to deeply assess the model generalization abilities as it has more or less the same size as a real-world screening task. Training, validation, and test sets were created using a ratio 80%:10%:10% for each target to minimize comparison bias.
2.1.2 Data set 2: ChEMBL_over
A second data set biased towards very unbalanced classes has been developed purposely for this work, that is ChEMBL_over. The data set was initially created through direct download using the ChEBML API. Specifically, the organism was set to human, and the relevant target_chembl_id and target_type for the protein were identified. Once these initial parameters were defined, the query was constructed by selecting the proper values of $IC_{50}$, $K_{d}$, $K_{i}$, $K_{dapp}$, Inhibition and Activity.
Each small molecule downloaded using the API was labeled based on the threshold values given in the literature. The molecules obtained in this way have a very low active/inactive ratio. For this reason, the ChEMBL_over data set was increased using an oversampling algorithm. This is a very common practice in machine learning to compensate for unbalanced data, and it has been adopted in recent years also in the chemical domain [8, 12]. We used a Random Over Sampling algorithm in the Imbalanced Learn library [19]. Overampling was applied on inactive molecules, until an active/inactive ratio of 1:15 was obtained. The generated molecules were carefully checked and validated through the RDKit library [18] to avoid introducing a classification bias due to the presence of chemically invalid molecules. Table 1 reports the amount of inactive molecules in the ChEMBL_over data set, arranged as “Active”, “I.C.” (Inactive Compound) and “I.O.” (Inactive Oversampled).
Protein Actives I.C. I.O.
ACK1 816 282 12240
ALK 2136 350 32040
CDK1 1310 769 19650
CDK2 2161 991 32385
CDK6 746 257 11190
CHK1 2475 684 22980
CK2A1 1145 174 18030
CLK2 786 263 99225
DYRK1A 1549 565 12390
EGFR 6947 2574 19755
ERK2 3735 721 37125
GSK3B 3120 1252 17175
INSR 1532 566 11790
IRAK4 2512 315 23235
ITK 1202 183 104205
JAK2 6615 667 56010
JNK3 826 758 46800
MAPK2K1 1342 421 37680
MELK 1317 290 20115
PDK1 1140 250 17100
Table 1: Summary of active and inactive molecules in the ChEMBL_over data set
2.2 Graph Convolutional Neural Network
The neural architecture employed in this study consists of two main components: a set of Graph Convolutional (GC) blocks and the classifier, which is implemented using linear layers. Each of the 20 single-target networks follows the same general architecture. The GC blocks are responsible for processing the graph-structured input data, capturing relationships between nodes through convolutional operations. Specifically, each block consists of a dedicated GC layer that implements the 1-dimensional Weisfeiler-Leman kernel (1-WL), as presented in [22], to enhance the network’s ability in extracting meaningful structural features. The dense layers serve as the final classifier, mapping the extracted features to the output predictions.
It is well known [11] that the global efficacy of a machine learning model for Virtual Screening is measured through the ability of correctly prioritizing the True Positives. Therefore, the main aim of the network is to provide a very high sensitivity across all the protein targets analyzed in this study. This focus on high sensitivity allows the network to deliver precise explanations of the molecular moieties involved in the target class assignment. To achieve this, we carefully optimized the network architecture, tailoring its structure to effectively balance complexity and performance.
The GC layers process the features in a U-Net-like fashion [27], initially expanding and then reducing the number of channels layer by layer, and ultimately resulting in a final representation with 32 features. The architecture consists of seven GC blocks, with the number of channels in each GC layer following this sequence: 32, 64, 128, 256, 128, 64, and again 32. Layer normalization and the ReLU activation function are used after each block.
Figure 1: The GCNN classifier architecture along with the HGE explainers
The convolutional layers described above use the message passing algorithm to stack data coming from the neighbors of each node in the graph. This aggregation process results in an updated set of node features placed along the channel dimension of the data sample. More formally, the algorithm uses a message function $M_{t}$ and a vertex update function $U_{t}$, which is usually implemented as a Multi-Layer Perceptron (MLP). Both $M_{t}$ and $U_{t}$ are learnable and differentiable functions to allow training in a neural network. Given the time-step $t$, the node $v$ collecting features, the edge $e_{vw}$ connecting node $v$ and any node $w$ in the set $N(v)$ of the neighbors of $v$ in the graph $G$, and the hidden states $h_{v}^{t}$ and $h_{w}^{t}$, the new message $m_{v}$ at time-step $t+1$ is calculated as:
(3) $m_{v}^{t+1}=\sum_{w\in N(v))}M_{t}(h_{v}^{t},h_{w}^{t},e_{vw})$
The new hidden state $h_{v}^{t+1}$ is then computed from $h_{v}^{t}$ and the new message $m_{v}^{t+1}$ using the vertex update function:
(4) $h_{v}^{t+1}=U_{t}(h_{v}^{t},m_{v}^{t+1})$
The augmented node feature vector coming from the GC operations flows into a Global Average Pooling (GAP) layer. This step is also known as the readout phase, and it represents a core operation when using the message passing algorithm. GAP condenses the information about all the nodes into a one-dimensional vector that averages all the features of the whole graph into a compressed and meaningful vector representation.
The second part of the network is a neural classifier consisting of four fully connected layers whose channels are, respectively, 32, 32, 16, and finally 1 that is fed to a sigmoid to estimate the probability for the input to be active on the target. ReLU activation functions and dropout layers are also added for each fully connected layer to add non-linearity correlation to the input data. The entire architecture of the single target network is reported in Figure 1 along with an example of the output of the explainers used in HGE.
A separate GCNN binary classifier was trained to assess compounds activity for each target. Weighted Binary Cross-Entropy was used with a higher weight value for the positives given the strong imbalance between active and inactive classes on targets. A grid search was performed to select the hyperparameters properly. The number of GC filters was searched in the set $\{1024,512,256,128,64,32,16\}$, while learning rates were in the range $[10^{-5},10^{-1}]$. The batch size ranged from $[8,64]$. The best model was stored after each epoch, and the early stopping callback was used to determine the ideal number of training epochs. The training process was conducted on an NVIDIA GeForce RTX 3090 with 10,496 CUDA cores, handling the extensive computations required for the grid search evaluation to determine the optimal hyperparameters.
2.3 Explainability analysis
As mentioned previously, the main contribution of this work is our Hierarchical Grad-CAM graph Explainer devised to gain information about the most relevant molecular moieties involved in protein-ligand binding prediction; different algorithms were tried in this respect. All selected XAI approaches belong to the so-called “instance-level methods”, that is, the ones that aim to identify the input features that contribute mainly to the prediction. At first, we applied GNNExplainer [37], a “perturbation-based method” that tracks the change in the output for minor variations of the same input data to determine which features are the most important for prediction. GNNExplainer provides information at a very fine-grained level, focusing on individual features. However, our goal was to obtain explanations at a higher level, specifically at the atom (i.e. node) level. As a consequence, we moved towards the class of “gradient and feature-based techniques”. An implementation from the work by Pope et al. [23] was utilized for this purpose. In this work, the authors implemented some graph-based counterparts of three well-established explainability methods, which had originally been designed for convolutional neural networks.
We chose Grad-CAM as the best performing technique for labeling pointwise contributions to positive predictions, as it provides valuable information on the most relevant features driving the model classification, particularly highlighting the key pharmacophoric elements within molecular graphs, as already mentioned in Section 1. This ability to visualize the model’s decision-making process, in addition to its suitability in the original image domain, where pixels are the data units and information flows between layers via the convolution operator, made it an ideal choice. In a GC layer, feature maps are built for each node of the graph by message passing (eq. 3). In this way, the features of each node encode the information flow from the node neighborhood, and the deeper the layer, the more its features collect information from many message passing steps computed in the previous layers (eq. 4). As a consequence, a feature map at a certain depth $d$ collects information from the $d-$neighbors of a node.
Grad-CAM aims at visualizing the most active region of the input when a particular class is predicted. This is done by examining the gradients of the target class score with respect to each feature map in the network layer under investigation. The weight $\alpha_{k}^{l,c}$, for class $c$ related to the $k$-th node feature at layer $l$ is computed as:
(5) $\alpha_{k}^{l,c}=\frac{1}{N}\sum_{n=1}^{N}\frac{\partial y^{c}}{\partial F_{k,n}^{l}}$
$F_{k,n}^{l}$ is the $k$-th feature value for the single node $n$ in layer $l$, while $y^{c}$ is the score for class $c$ obtained from the last network layer, before the output unit. Obviously, the partial derivative term is computed using the chain rule when $l$ is not the last layer. Grad-CAM values for the entire node $n$ are summed over its features, and passed to the ReLU activation function to filter out negative contributions to the prediction of class $c$.
(6) $L_{Grad-CAM}[l,n]=\operatorname{ReLU}\left(\sum_{k}\alpha_{k}^{l,c}F_{k,n}^{l}\right)$
In our implementation, we set $c=1$, that is, we have a single class score, as the explainer was developed for a binary classifier. Moreover, the ReLU activation function has been removed to capture not only positive contributions to the prediction, but also negative ones. Our choice is motivated by the intrinsic nature of the classification task to be explained. In the image domain, classification tasks can involve assigning each pixel to a specific class, as in semantic segmentation, or labeling the entire image based on the presence of class-discriminative features learned by the network. In both cases, classification can be interpreted as a pattern recognition task. Consequently, the explainer is designed to highlight “regions” that contribute to the prediction while disregarding irrelevant areas, which is the intended purpose of the ReLU activation function.
In this context, bioactivity prediction is a classification task in which molecules are labeled positive or negative if they “globally” trigger a particular biological event or not. Many factors are involved in bioactivity: the presence of suitable pharmacophoric moieties, molecular chirality, solubility, etc. All of them can be captured by training the GCNN on a three-dimensional molecular representation with rich information at the nodes, but cannot be directly related to just “positive” or “negative” Grad-CAM values at each node. As a consequence, we removed ReLU for our explainability task. Moreover, pharmacophoric information is strictly related to the atom connectivity within the molecular structure.
In view of this last consideration, we evaluated Grad-CAM multiple times at different depths in the convolutional segment of the network. We used the classical Grad-CAM arrangement after the last layer of the network, which is the seventh. This explainer aims to highlight relevant features globally and is unable to provide information about subtle chemical and pharmacophoric moieties like H-bond acceptors or donors. To reach this aim, we placed also a Grad-CAM explainer at the very first GC layer to consider the atoms’ impact in a 1-neighborhood that is at a bond level. Another explainer was placed after the third layer to catch information related to 3-neighborhoods. At this level, as we analyze the 3-neighborhood of each atom, the action of an entire substructure like aromatic ring can be explained.
The whole explanation map has been built as follows. At first, the three Grad-CAM value distributions have been summed for each node:
(7) $L_{Grad-CAM}[n]=L_{Grad-CAM}[1,n]+L_{Grad-CAM}[3,n]+L_{Grad-CAM}[7,n]$
The resulting $L_{Grad-CAM}$ distribution has been min-max scaled, and only contributions higher than $0.7$ have been regarded as relevant. We called the final output “merged output”. It maps all of the key chemical features collected from the various explainer layers, combining them to create a global view of each of these levels’ contributions. This approach will determine the whole chemical interacting feature assignment. The GCN architectures are in fact able to recognize molecular graphs and classify the molecules based on their nodes and connections. However, atomic typing and identification of the most important chemical groups in determining the assignment of a molecule to a specific target are lacking. In order to account this issue, we employed HGE explanations to contextualise the information obtained. To achieve this, it was necessary to harness the power of the software package RDKithttps://www.rdkit.org/. Actually, the software has the capability to identify pharmacophoric regions for each of the molecular sub-structures found in the molecular structure. Once the moieties are labeled, they are cross-referenced with the results obtained from HGE to select only those that contributed to the bioactivity classification. This procedure is crucial for each layer of the explainer and allows us to chemically translate the information obtained from the various layers of the GCN. Looking at Figure 1, H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green. Lastly, moieties without an explicit pharmacophoric label where $L_{Grad-CAM}>0.7$ are labeled as “relevant”, and are colored in red. In Section 3.4 a detailed analysis is reported about the structures highlighted by HGE.
3 Results
Experiments were arranged in three stages. At first, a GCNN classifier was trained for each target, to allow HGE computation. Furthermore, the explainability analysis was performed on 143 known kinase inhibitors from DrugBankhttps://go.drugbank.com/), and about 80% of them were correctly assigned to the annotated target. Finally, our GCNN architectures were compared with Chemprop, one of the most widely used frameworks for training message-passing neural networks aimed at predicting molecular properties.
3.1 Metrics
To test the predictive capabilities of the proposed architectures, we used two families of metrics. At first we reported the classical Machine Learning scores used for classification. The Balanced Accuracy assesses the network’s ability to make correct predictions despite the class imbalance, while the Sensitivity (also named Recall) focuses solely on the network’s performance in correctly classifying active molecules. The F1-score quantifies how well the model can correctly identify positive instances (Recall) while avoiding false positives (Precision). On the other hand, the Area Under the Curve (AUC) provides a global measure of the model’s ability to discriminate between positive and negative instances, and it is computed from the Receiver Operating Characteristics (ROC) curve that plots the Sensitivity vs. the False Positive Rate varying the threshold to discriminate positives and negatives.
Prioritizing the best active ligands is the true outcome of a VS task. For this purpose, the True Positive vs Positives ratio (TP/P) and the Enrichment Factor (EF) are used. TP/P is employed to assess the correct predictions count relative to the total positives in a given class, the EF metric measures how many times the tested model outperforms a purely random process in predicting a given class. Both measures are computed for a given topmost percentage of the test set. We used, respectively, 1%, 2%, 5%, and 10% of the whole test set, according to the chemical literature.
3.2 Results on EMBER data set
The results on the EMBER data set for each single-target classifier are shown in Table 2 and Table 3. Table 2 shows the results achieved by our architecture for each target, and the last two columns show the sensitivity comparison with Chemprop. The comparison with Chemprop is reported in Table 4 where all metrics averaged on the 20 targets are reported for both approaches.
As it can be seen, F1-Score values show that the network is able to overcome the data set imbalance while simultaneously reducing the number of false negatives and false positives. The use of a properly balanced loss and a soft F-score worked fine for the classification task. Particularly noticeable is the Sensitivity value that is higher than that of Chemprop for all the targets.
Table 2: Performance metrics of our GCNN architecture on EMBER data set.
Protein Bal. Accuracy F1-Score AUC G-Sensitivity C-Sensitivity
ACK 0.891 0.979 0.983 0.811 0.657
ALK 0.929 0.972 0.987 0.893 0.742
CDK1 0.921 0.956 0.969 0.907 0.650
CDK2 0.932 0.940 0.971 0.950 0.721
CDK6 0.916 0.989 0.988 0.846 0.772
CHK1 0.923 0.954 0.977 0.912 0.604
CK2A1 0.942 0.979 0.989 0.911 0.705
CLK2 0.938 0.929 0.989 0.965 0.881
DYRK1A 0.913 0.983 0.986 0.847 0.730
EGFR 0.930 0.982 0.990 0.881 0.749
ERK2 0.927 0.941 0.977 0.935 0.782
GSK3B 0.903 0.970 0.978 0.847 0.615
INSR 0.893 0.974 0.977 0.823 0.430
IRAK4 0.930 0.953 0.979 0.931 0.619
ITK 0.928 0.921 0.981 0.958 0.814
JKA2 0.945 0.945 0.984 0.960 0.786
JNK 0.918 0.923 0.971 0.946 0.691
MAP2K1 0.944 0.969 0.991 0.926 0.807
MELK 0.916 0.969 0.982 0.874 0.635
PDK1 0.913 0.973 0.986 0.860 0.693
mean 0.899 0.704
*G-Sensitivity = Our GCN Sensitivity; C-Sensitivity = Chemprop Sensitivity
Protein TP/P 1% TP/P 2% TP/P 5% TP/P 10% EF 1% EF 2% EF 5% EF 10%
ACK 72/106 78/106 94/106 100/106 67 36 17 9
ALK 123/254 186/254 227/254 242/254 48 36 17 9
CDK1 80/205 117/205 177/205 190/205 39 28 17 9
CDK2 86/303 141/303 240/303 287/303 28 23 15 9
CDK6 82/104 89/104 94/104 99/104 78 42 18 9
INSR 89/217 123/217 180/217 204/217 41 28 16 9
ITK 98/158 127/158 145/158 154/158 62 40 18 9
JAK2 134/832 268/832 651/832 774/832 16 16 15 9
JNK3 79/105 83/105 93/105 102/105 75 39 17 9
MELK 125/185 158/185 171/185 179/185 67 42 18 9
CHK1 126/343 189/343 269/343 319/343 36 27 15 9
CK2A1 93/151 115/151 128/151 141/151 61 38 16 9
CLK2 52/102 63/102 85/102 96/102 50 30 16 9
DYRK1A 72/174 106/174 144/174 167/174 41 30 16 9
EGFR 129/702 262/702 520/702 641/702 18 18 14 9
ERK2 130/525 260/525 444/525 503/525 24 24 16 9
GSK3 102/393 174/393 293/393 358/393 25 22 14 9
IRAK4 134/339 249/339 311/339 327/339 39 36 18 9
MAP2K1 106/191 133/191 166/191 181/191 55 34 17 9
PDK1 108/187 132/187 164/187 183/187 57 35 17 9
Table 3: TP/P and EF results for the EMBER data set
Model Bal. Accuracy Sensitivity F1-Score AUC
Chemprop 0.851 0.704 0.756 0.851
Our GCNN 0.923 0.899 0.960 0.982
Table 4: Average evaluation metrics for Chemprop and our GCNN on EMBER data set
The high Sensitivity indicates a good ability of our GCNNs to correctly prioritize active ligands, and this is confirmed by the optimal performance obtained in terms of TP/P and EF values.
3.3 Results on ChEMBL_over data set
We carried out the same experiments on the ChEMBL_over data set as on the EMBER one. The results obtained from the 20 GCNNs trained purposely are reported in table 5, table 6, and table 7.
Table 5: Performance metrics of our GCNN architecture on ChEMBL_over data set.
Protein Bal. accuracy AUC F1-score G-Sensitivity* C-Sensitivity*
ACK1 0.972 0.992 0.954 0.947 0.866
ALK 0.979 0.996 0.978 0.957 0.944
CDK1 0.949 0.995 0.933 0.901 0.939
CDK2 0.953 0.985 0.862 0.929 0.907
CDK6 0.952 0.951 0.901 0.914 0.893
CHK1 0.969 0.991 0.965 0.939 0.907
CK2A1 0.986 1.000 0.986 0.972 0.896
CLK2 0.956 0.991 0.954 0.911 0.899
DYRK1A 0.943 0.974 0.871 0.903 0.723
EGFR 0.914 0.961 0.715 0.887 0.895
ERK2 0.954 0.979 0.803 0.950 0.893
GSK3B 0.958 0.984 0.876 0.937 0.869
INSR 0.981 0.990 0.981 0.962 0.928
IRAK4 0.993 0.999 0.993 0.986 0.952
ITK 0.997 1.000 0.997 0.995 0.950
JAK2 0.993 0.998 0.980 0.989 0.970
JNK3 0.942 0.979 0.707 0.959 0.819
MAPK2K1 0.964 0.989 0.963 0.928 0.903
MELK 0.992 0.999 0.992 0.984 0.833
PDK1 0.988 1.000 0.988 0.977 0.947
mean 0.946 0.897
*G-Sensitivity = Our GCN Sensitivity; C-Sensitivity = Chemprop Sensitivity
Protein TP/P 1% TP/P 2% TP/P 5% TP/P 10% EF 1% EF 2% EF 5% EF 10%
ACK1 9/76 17/76 44/76 73/76 12 11 12 10
ALK 23/211 46/211 115/211 202/211 11 11 11 10
CDK1 14/131 29/131 72/131 125/131 11 11 11 10
CDK2 25/225 49/225 123/225 208/225 11 11 11 9
CDK6 8/70 16/70 39/70 64/70 11 11 11 9
CHK1 28/263 57/263 142/263 248/263 11 11 11 9
CK2A1 12/108 25/108 62/108 108/108 11 12 11 10
CLK2 9/79 18/79 44/79 75/79 11 11 11 10
DYRK1A 17/154 34/154 84/154 140/154 11 11 11 9
EGFR 71/655 143/655 347/655 539/655 11 11 11 8
ERK2 40/361 79/361 198/361 332/361 11 11 11 9
GSK3B 54/491 109/491 271/491 460/491 11 11 11 9
INSR 28/262 57/262 142/262 257/262 11 11 11 10
IRAK4 41/368 81/368 204/368 365/368 11 11 11 10
ITK 21/185 41/185 104/185 185/185 11 11 11 10
JAK2 112/1040 225/1040 562/1040 1031/1040 11 11 11 10
JNK3 13/122 27/122 66/122 110/122 11 11 11 9
MAPK2K1 22/194 45/194 111/194 189/194 11 12 11 10
MELK 38/364 76/364 190/364 361/364 10 10 10 10
PDK1 20/173 39/173 98/173 173/173 12 11 11 10
Table 6: TP/P and EF results for the ChEMBL_over data set
Model Bal. Accuracy Sensitivity F1-Score AUC
Chemprop 0.948 0.897 0.945 0.982
Our GCNN 0.967 0.946 0.920 0.988
Table 7: Average evaluation metrics for Chemprop and our GCNN on ChEMBL_over data set.
The performance of our GCNNs on the ChEMLB_over data set remains consistent with the one obtained on the EMBER one. Our architectures performs better than Chemprop on 18 out 20 targets in terms of Sensitivity, but also in the target when they perform less than Chemprop the difference is very subtle. The TP/P and EF scores are high on all the targets. Finally, Chemprop reports an average F1-score that is higher than our architectures. This is not a limitation because our architectures were trained for reaching high Sensitivity as it is demonstrated by the average of such a metric on all the target. As a consequence, our networks perform better than Chemprop in prioritizing True Positive. This result could reveal a vulnerability in the Chemprop net, which seems more prone than our GCNNs to be misled in assessing True Positives by the presence of over-sampled molecules.
3.4 Explainability results
To evaluate the explainability power of HGE, we selected a test set made by 143 known drugs from DrugBank https://go.drugbank.com/) for the 20 kinases considered in this work to compare the predictions with experimental and clinical results. For this purpose, only drugs flagged as approved and inhibitor on DrugBank were chosen. For each drug chosen, correct target assignation was evaluated and the relevant chemical moieties were collected and analysed. For each chemotype and for each target, the consistency of key chemical features identification was assessed. Thus, 116 drugs out of 143 were correctly assigned to the target known in the literature. In order to further stress the discriminative capacity of the proposed approach, 13 drugs with bioactivity on multiple targets have been identified. The results obtained from the evaluation phase of the results are shown in table 8. Finally, to further stress the capabilities of the proposed explainer, 13 drugs with multi-target bioactivity were tested to see if the explainer’s predictions were kept consistent among different targets annotated. The results is shown in table 8.
Target # of Drugs *C.P.D.
ACK 1 0
CDK2 5 4
CDK6 3 3
CHK1 6 4
C2KA1 4 4
CLK2 1 1
DYRK1A 1 1
ERK2 5 1
GSK3b 7 4
INSR 1 1
IRAK4 1 0
ITK 2 2
JAK2 2 2
JNK3 1 1
MAPK2K1 2 1
MELK 1 0
PDK1 1 1
*C.P.D. = Correctly Predicted Drugs
Table 8: Classification and explainability predictions on the test drugs
As shown in the table 8, only 30/44 (0.6818) drugs with multi-target activity are correctly classified, but this result can be interpreted positively. In fact, the same molecules were tested with HGE, demonstrating the robustness in correctly classifying molecular scaffolds. All the drugs have been tested with HGE, in order to assess whether the core pharmacophore could be predicted correctly, and these results are reported in supplementary material 3. Some examples are shown in figure 2, where the prediction of the molecular moieties of Apigenin, a drug that is bioactive on CDK6 and CK2A1, remains congruous, despite the bioactivity is calculated on two different targets. This result demonstrates that HGE is consistent and robust when discriminating between compounds even when the neural network is trained on different protein targets.
Figure 2: Predictions obtained with HGE for Apigenin on CDK6 and CK2A1
Figure 3 shows the prediction of HGE for Trilaciclib as another example of this consistent behavior. This drug is active on both CDK2 and CDK6 whose binding sites are very similar. This evidence remains congruous for all the other 11 drugs (see supplementary material S3), demonstrating how HGE manages to maintain its predictive capacity of molecular moieties.
Figure 3: Predictions obtained with HGE for Trilaciclib on CDK2 and CDK6
Considering the explainer’s ability to correctly identify the moieties for drugs with multi-target activity, it was been decided to further test the robustness of HGE. For this phase, different drugs that are active on the same target have been identified whose molecular scaffold was chemically similar. Some examples are shown in figure 4 which highlights the predicted moieties in Apigenin and Chrysin (both active on CDK6) and in Pacritinib and SB1578 (both active on JAK2).
Figure 4: Chemical features identified by HGE in a) Apigenin and Chrysin when assigned to CDK6 protein and b) Chemical features identified by HGE in Pacritinib and SB1578 when assigned to JAK2 protein
3.5 Discussion
Experimental results show clearly that HGE performs very well in real applications with respect to both the prediction task and the explanation of the molecular moieties. The main strength of the presented approach is the combination of different explanation layers (from single atoms to complex structures). Each layer is capable of catching different aspects of the molecule under investigation, and each step enriches the others, giving a complete interpretation of the molecular moieties useful to be classified as active for a specific target. As demonstrated in Figures 2, 3 and 4, the moieties are predicted consistently across similar chemotypes for the annotated target, and the HGE prediction is consistent and robust even for drugs with bioactivity on distinct targets.
4 Conclusion
In this work the HGE framework has been proposed that combines different Grad-CAM explainability layers arranged in a message passing hierarchy to detect the molecular moieties involved in predicting the activity of a compound versus a given protein target. To achieve our result, 20 GCNN classifiers have been trained on two different data sets derived from CheMBL to predict ligands’ activity for 20 Kinase proteins. Our approach proved to be very effective in classification and was compared with Chemprop in this respect. Moreover, the moieties explanation ability was assessed using a suitable test set of 143 approved drugs derived from DrugBank, which are active on the 20 targets and comparing them with a proper docking analysis carried out by expert computational chemists.
Our experiments produced SOTA results with respect to classification and the moieties detection is highly precise. Such results set the framework as a very useful tool for Virtual Screening where our approach unveils potential hidden molecular moieties useful for binding a specific target. Finally, HGE is able to associate a given set of moieties with a particular target. As a consequence, very interesting insights can also be gained in the direction of drug repurposing if we use HGE to analyze a known drug that is known to own some of the moieties related to such a target. HGE performance is shown to be good and reliable both in predicting the same moieties for a target (figure 4) and in identifying the same molecular features of the same molecule on different targets (figure 2, 3). These two complementary aspects demonstrate, in our opinion, a good reliability of HGE in identifying the important chemical moieties for small molecule-target recognition. Future research will be dedicated to studying the ability of HGE to re-propose drugs and to understanding the link between classification properties and explainer prediction for assessing how to increase the amount of information that can be obtained from the explainer.
5 Data and Code Availability
6 Supplementary information
Supplementary file S1 gives a detailed description of the data used to construct the EMBER data set. In addition, there is a table showing the results obtained from the test with drugs active on multiple targets. Supplementary file S2 contains the results obtained from the HGE prediction for all the selected drugs, while supplementary file S3 contains the HGE plots of the drugs with activity on multiple targets reported to compare their robustness in the prediction of the molecular scaffold.
Abbrevations
  • CADD. Computer-Aided Drug Design
  • VS. Virtual Screening
  • SAR. Structure Activity Relationship
  • XAI. Explainable AI
  • EMBER. EMBedding multiplE molecular fingeRprints
  • HGE. Hierarchical Graph Explainer
  • IC50. Half Maximal Inhibitory Concentration
  • I.C. Inactive Compound
  • I.O. Inactive Oversampled
  • GC Graph Convolution
  • GCNN Graph Convolution Neural Network
  • GAP Global Average Pooling
  • Ki. Inhibition constant
  • Kd. Dissociation constant
  • MLP. Multi Layer Perceptron
  • PDB. Protein Data Bank
  • ReLU. Rectified Linear Unit
  • wBCE. Weighted Binary Cross-Entropy
  • CAM Class Activation Map
  • Grad-CAM. Gradient-weighted Class Activaction Mapping
  • TP. True Positive
  • P. Positive
  • TN. True Negative
  • N. Negative
  • TP/P. True Positive vs Positive ratio
  • AUC Area Under the Curve
  • EF. Enrichment Factor
7 Acknowledgments
This work has been partially supported by project:
1.“SAMOTHRACE” (Sicilian MicronanoTech Research And Innovation Center), cup project B73C22000810001, project code ECS_00000022.
2. “DARE” (DigitAl lifelong pRevEntion), cup project B53C22006450001, project code PNC_0000002.